OpenStudy (anonymous):

An object weighs 100 lb. on the surface of the earth. How much does it weigh 4R from the surface? (5R from the center)

OpenStudy (ujjwal):

The weight of an object is nothing else than the gravitational force of attraction. Now, use the fact that Gravitational force is inversely proportional to square of distance between the object and center of earth and solve it!

OpenStudy (anonymous):

So it would be force = mass / R^2 ? Giving you force = 100 / 25 = 4. Right ? & Thanks for helping !

OpenStudy (ujjwal):

No, it won't be like that! You will have the relation \[\frac{F_1}{F_2}=\frac{r_2^2}{r_1^2}\]where, \(F_1\)=100 \(r_1\)=R \(r_2\)=5R You are now supposed to find \(F_2\)

OpenStudy (anonymous):

Oh, I just saw this. I am so sorry for taking so long to reply ! & Ohh. That would be why I missed it on my homework :P Haha. But it's okay, thanks for helping !